OpenStudy (anonymous):

MEDALS AND FAN 1 Find the exact area of a circle having the given circumference. 4(3) A = 2√3 4√3 12 2 Find the exact area of the region bounded by two concentric circles with radii 10 inches and 6 inches. 16 64 88 3 Find the area of a circle whose area equals the half of the sum of the areas of two circles with radii 2 and 3. 13 6.5 5

OpenStudy (aric200):

One question Per post I will do #1 1. Find the exact area of a circle having the given circumference. 4(3) A = A. 2√3 B. 4√3 C. 12 Circumference = 2r(pi) = 4(sqrt3pi) r(pi) = 2(sqrt3pi) r = 2(sqrt3pi)/pi Now we can find area ------------------ area = pi(r^2) = pi(2(sqrt3pi)/pi)^2 = pi(12pi/pi^2) = 12pi^2/pi^2 So the answer is C. 12